Find the value of (2 1/3 * 7/8 ) - ( 5/6 * 1 2/5)
if x * y = (x+y)/(x-Y)

Answer by palanisamy(496)

Given, x*y = (x+y)/(x-y)
2 1/3 * 7/8 = 7/3 * 7/8
2 1/3 * 7/8 = (7/3+7/8)/(7/3-7/8)
2 1/3 * 7/8 = [(56+21)/24]/[(56-21)/24]
2 1/3 * 7/8 = 77/35
2 1/3 * 7/8 = 11/5 ...(1)
Next,
5/6* 1 2/5 = 5/6 *7/5
5/6* 1 2/5 = (5/6 +7/5)/(5/6-7/5)
5/6* 1 2/5 = [(25+42)/30]/[(25-42)/30]
5/6* 1 2/5 = 67/(-17)
5/6* 1 2/5 = - 67/17 ...(2)
Now,
(1)-(2):
2 1/3*7/8 - 5/6 * 1 2/5 = 11/5-(-67/17)
= (187+335)/85
= 522/85